OpenStudy (anonymous):

A cookie recipe uses 1/2 teaspoon of vanilla with 3/4 cup of flour. how much vanilla should be used with five cups of flour

OpenStudy (anonymous):

what is \(5\div \frac{3}{4}\)

OpenStudy (anonymous):

if you do not know, that is fine, i will tell you just asking

OpenStudy (anonymous):

6.6...?

OpenStudy (anonymous):

fractions, not decimals

OpenStudy (anonymous):

\[5\div \frac{3}{4}=5\times \frac{4}{3}=\frac{20}{3}\]

OpenStudy (anonymous):

which is \(6\tfrac{2}{3}\) like you had, but we need the fraction now that we know there are \(\frac{20}{3}\) three quarter cups in five cups, and each one take half a teaspoon, we can multiply \[\frac{20}{3}\times \frac{1}{2}=\frac{10}{3}\]

OpenStudy (anonymous):

so \(3\tfrac{1}{3}\) teaspoons per for five cups

OpenStudy (anonymous):

thanks for going step by step!